# APOLLO: Automated LLM and Lean Collaboration for Advanced Formal Reasoning

### v2

@@ Abstract
Formal reasoning and automated theorem proving constitute a challenging subfield of machine learning, in which machines are tasked with proving mathematical theorems using formal languages like Lean. A formal verification system can check whether a formal proof is correct or not almost instantaneously, but generating a completely correct formal proof with large language models (LLMs) remains a formidable task. The usual approach in the literature is to prompt the LLM many times (up to several thousands) until one of the generated proofs passes the verification system. In this work, we present APOLLO (Automated PrOof repair via LLM and Lean cOllaboration), a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with an LLM’s reasoning abilities to achieve better proof-generation results at a low sampling budget. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The repaired sub-proofs are recombined and reverified, iterating up to a user-controlled maximum number of attempts. On the miniF2F benchmark, we establish a new state-of-the-art accuracy of 75.0% among 7B-parameter models while keeping the sampling budget below one thousand. Moreover, Apollo raises the state-of-the-art accuracy for Goedel-Prover-SFT to 65.6% while cutting sample complexity from 25,600 to a few hundred. General-purpose models (o3-mini, o4-mini) jump from 3–7% to over 40% accuracy. Our results demonstrate that targeted, compiler-guided repair of LLM outputs yields dramatic gains in both efficiency and correctness, suggesting a general paradigm for scalable automated theorem proving.
@@ 3 Our Approach
Our compiler of choice is the Lean REPL [38], a “Read–Eval–Print–Loop” for Lean4. The REPL provides a complete list of compilation errors and warnings, each with a detailed description and the location (for example, the line number) where it occurred. Errors can include syntax mistakes, nonexistent tactics, unused tactics, and more. We aggregate this information to drive our agent’s dynamic code repair and sub-problem extraction.
@@ 4 Experiments
In this section, we present empirical results for Apollo on the MiniF2F dataset [8], which comprises of 488 formalized problems drawn from AIME, IMO, and AMC competitions. The benchmark is evenly split into validation and test sets (244 problems each); here we report results on the MiniF2F-test partition. To demonstrate Apollo’s effectiveness, we evaluate it on a range of state-of-the-art whole-proof generation models. All experiments use Lean v4.17.0 and run on eight NVIDIA A5000 GPUs with 128 CPU cores. We used @$32$ sampling during sub-proof generation unless stated otherwise. Baseline numbers are sourced from each model’s original publication works.
@@ 4.1 The effect of applying Apollo on top of the base models
Figure 3 shows the impact of Apollo on two state-of-the-art whole-proof generation models: Goedel Prover-SFT [31] and Kimina-Prover-Preview-Distill-7B [34]. Applying Apollo increases Goedel-Prover-SFT’s top accuracy from 64.7% to 65.6% while reducing its sampling budget by two orders of magnitude (from 25,600 generated samples to only a few hundred on average). Similarly, Kimina-Prover-Preview-Distill-7B achieves a new state-of-the-art accuracy of 75.0% with roughly one-third of the previous sample budget. We report the exact numbers in Table 1.
@@
Figure 3: Performance of base models against the Apollo aided models on MiniF2F-test dataset at different sampling budgets ranging from 32 up to 25,600.
Table 1: Sampling cost and accuracy before vs. after applying Apollo. Cost refers to the sampling budget parameter $K$ in whole-proof generation models and average number of generated samples for Apollo. The best accuracy results are bolded.
@@
Cost Accuracy (%) Cost Accuracy (%) Cost Accuracy (%) Cost Accuracy (%)
32 57.6 32 57.6 32 63.1 32 63.1
64 59.2 80 60.7 96 68.9
3200 62.7 151 63.5 191 74.2
25,600 64.7 362 65.6 1024 70.8 307 75.0
@@ 4.2 Comparison with SoTA LLMs
Table 2 compares whole-proof generation and tree-search methods on MiniF2F-test. For each approach, we report its sampling budget, defined as the top-$K$ parameter for LLM generators or equivalent search-depth parameters for tree-search provers. Since Apollo’s budget varies with recursion depth $r$, we instead report the mean number of proof sampled during procedure.
When Apollo is applied on top of any generator, it establishes a new best accuracy for that model. For instance, Goedel-Prover-SFT’s accuracy jumps to 65.6%, surpassing the previous best result (which required 25,600 sample budget). In contrast, Apollo requires only 362 samples on average. Likewise, Kimina-Prover-Preview-Distill-7B sees a 4.2% accuracy gain while reducing its sample budget below the original 1,024. Overall, Apollo not only boosts accuracy but does so with a smaller sampling budget, setting a new state-of-the-art result among sub-7B-parameter LLMs.
@@
Table 2: Comparison of Apollo accuracy against state-of-the-art models on the miniF2F-test dataset.
Method Model size Sample budget miniF2F-test
@@
Hypertree Proof Search [26] 600M $64\times 5000$ $41.0\%$
InternLM2.5-StepProver + BFS + CG [24] 7B $256\times 32\times 600$ $65.9\%$
HunyuanProver v16 + BFS + DC [41] 7B $600\times 8\times 400$ $68.4\%$
BFS-Prover [25] 7B $2048\times 2\times 600$ $70.8\%$
@@
DeepSeek-R1-Distill-Qwen-7B [42] 7B 32 42.6%
Leanabell-Prover-GD-RL [32] 7B $128$ $61.1\%$
STP [33] 7B $25600$ $67.6\%$
o3-mini [39] - 1 $3.3\%$
$32$ $24.6\%$
o4-mini [40] - 1 $7.0\%$
Goedel-Prover-SFT [31] 7B 32 $57.6\%$
$3200$ $62.7\%$
$25600$ $64.7\%$
Kimina-Prover-Preview-Distill [34] 7B 1 $52.5\%$
$32$ $63.1\%$
$1024$ $70.8\%$
@@
o3-mini + Apollo - 8 40.2% (+33.2%)
o4-mini + Apollo - 15 46.7% (+39.7%)
Goedel-Prover-SFT + Apollo 7B 362 65.6% (+0.9%)
Kimina-Prover-Preview-Distill-7B + Apollo 7B 307 75.0% (+4.2%)
@@ 4.3 Distribution of Proof Lengths
Figure 4: Distribution of proof lengths for selected models before (left) and after (right) applying Apollo. In the “before” plots, proof lengths cover only proofs the base model discharged without the help of Apollo; in the “after” plots, proof lengths cover only proofs completed with Apollo’s assistance.
@@ 4.4 Impact of recursion depth $r$ on proof accuracy
We evaluate how the recursion-depth parameter $r$ affects accuracy on MiniF2F-test. A recursion depth of $r=0$ corresponds to the base model’s standalone performance; higher $r$ values permit additional rounds of Apollo. Figure 5 plots accuracy versus $r$ for three models: o4-mini ($r=0\ldots 4$), Kimina-Prover-Preview-Distill-7B ($r=0\ldots 5$), and Goedel-Prover-SFT ($r=0\ldots 5$).
@@ 6 Conclusion
In this work, we present Apollo, a novel, modular pipeline that combines syntax cleaning, automatic solvers, and LLM-driven sub-proof generation to transform an LLM’s initial proof sketch into a fully verified Lean4 proof. This framework harnesses the full power of the Lean compiler and integrated solvers, merging them with LLM systems. Applied across four whole-proof generation models, ranging from general-purpose LLMs (o3-mini, o4-mini) to specialized provers (Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT), Apollo consistently establishes new best accuracies on the miniF2F benchmark while reducing sampling budgets by one to two orders of magnitude. Our empirical analysis shows that Apollo not only raises overall proof success rates but also guides models to generate longer, more structured proofs, as evidenced by a drastic increase in average proof length. We further demonstrate how the recursion-depth parameter $r$ trades off sample complexity against accuracy, achieving robust gains with only a few hundred samples. We believe that Apollo’s collaboration between LLMs, automatic solvers, and the Lean compiler defines a paradigm of agentic systems that produce high-quality, verifiable proofs for increasingly challenging theorems.

### v3

@@ Abstract
Formal reasoning and automated theorem proving constitute a challenging subfield of machine learning, in which machines are tasked with proving mathematical theorems using formal languages like Lean. A formal verification system can check whether a formal proof is correct or not almost instantaneously, but generating a completely correct formal proof with large language models (LLMs) remains a formidable task. The usual approach in the literature is to prompt the LLM many times (up to several thousands) until one of the generated proofs passes the verification system. In this work, we present APOLLO (Automated PrOof repair via LLM and Lean cOllaboration), a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with an LLM’s reasoning abilities to achieve better proof-generation results at a low sampling budget. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The repaired sub-proofs are recombined and reverified, iterating up to a user-controlled maximum number of attempts. On the miniF2F benchmark, we establish a new state-of-the-art accuracy of 84.9% among sub 8B-parameter models while keeping the sampling budget below one hundred. Moreover, Apollo raises the state-of-the-art accuracy for Goedel-Prover-SFT to 65.6% while cutting sample complexity from 25,600 to a few hundred. General-purpose models (o3-mini, o4-mini) jump from 3–7% to over 40% accuracy. Our results demonstrate that targeted, compiler-guided repair of LLM outputs yields dramatic gains in both efficiency and correctness, suggesting a general paradigm for scalable automated theorem proving. The codebase is available at https://github.com/aziksh-ospanov/APOLLO
@@ 3 Our Approach
Our compiler of choice is the Lean REPL [38], a “Read–Eval–Print–Loop” for Lean4. The REPL provides a complete list of compilation errors and warnings, each with a detailed description and the location (for example, the line number) where it occurred. Errors can include syntax mistakes, nonexistent tactics, unused tactics, and more. We aggregate this information to drive our agent’s dynamic code repair and sub-problem extraction. The pseudo-code for our framework can be found in Algorithm 1 in the Appendix.
@@ 4 Experiments
In this section, we present empirical results for Apollo on the MiniF2F dataset [8], which comprises of 488 formalized problems drawn from AIME, IMO, and AMC competitions. The benchmark is evenly split into validation and test sets (244 problems each); here we report results on the MiniF2F-test partition. To demonstrate Apollo’s effectiveness, we evaluate it on a range of state-of-the-art whole-proof generation models. All experiments use Lean v4.17.0 and run on eight NVIDIA A5000 GPUs with 128 CPU cores. We used @32@32 sampling during sub-proof generation unless stated otherwise. Baseline numbers are sourced from each model’s original publication works.
@@ 4.1 The effect of applying Apollo on top of the base models
Figure 3 shows the impact of Apollo on two state-of-the-art whole-proof generation models: Goedel Prover-SFT [31] and Kimina-Prover-Preview-Distill-7B [34]. Applying Apollo increases Goedel-Prover-SFT’s top accuracy from 64.7% to 65.6% while reducing its sampling budget by two orders of magnitude (from 25,600 generated samples to only a few hundred on average). Similarly, Kimina-Prover-Preview-Distill-7B achieves a new best Kimina-Prover-Preview-Distill-7B accuracy of 75.0% with roughly one-third of the previous sample budget. We report the exact numbers in Table 1.
(a) Model accuracy against sampling budget
(b) Model accuracy against generated token budget
Figure 3: Performance of base models against the Apollo aided models on MiniF2F-test dataset at different sample budgets and token length.
@@
Table 1: Table results comparing sampling cost, accuracy, and token usage before and after applying Apollo. “Cost” denotes the sampling budget $K$ for whole-proof generation and the recursion depth $r$ for Apollo. Column NN reports the average number of tokens generated per problem in "Chain-of-Thought" mode. Bold cells highlight the best accuracy. Results are shown on the miniF2F-test dataset for two models: Kimina-Prover-Preview-Distill-7B and Goedel-SFT.
@@
$K$ NN Acc. (%) $r$ NN Acc. (%) $K$ NN Acc. (%) $r$ NN Acc. (%)
32 16.1K 57.6% 0 16.1K 57.6% 32 140K 63.1% 0 140K 63.1%
64 31.8K 59.2% 2 38.3K 60.7% 2 406K 68.9%
3200 1.6M 62.7% 4 74.6K 63.5% 4 816K 74.2%
25600 12.7M 64.7% 6 179.0K 65.6% 1024 4.5M 70.8% 6 1.3M 75.0%
@@ 4.2 Comparison with SoTA LLMs
Table 2 compares whole-proof generation and tree-search methods on MiniF2F-test. For each approach, we report its sampling budget, defined as the top-$K$ parameter for LLM generators or equivalent search-depth parameters for tree-search provers. Since Apollo’s budget varies with recursion depth $r$, we instead report the mean number of proof sampled during procedure. Moreover, we report an average number of generated tokens as another metric for computational cost of proof generation. For some models due to inability to collect generated tokens, we leave them blank and report sampling budget only.
When Apollo is applied on top of any generator, it establishes a new best accuracy for that model. For instance, Goedel-Prover-SFT’s accuracy jumps to 65.6%, surpassing the previous best result (which required 25,600 sample budget). In contrast, Apollo requires only 362 samples on average. Likewise, Kimina-Prover-Preview-Distill-7B sees a 4.2% accuracy gain while reducing its sample budget below the original 1,024. To further validate Apollo’s generalizability, we tested the repair framework on Goedel-V2-8B [41], the current state-of-the-art theorem prover. We observe that, at similar sample and token budgets, Apollo achieves a 0.9% accuracy gain, whereas the base LLM requires twice the budget to reach the same accuracy. Overall, Apollo not only boosts accuracy but does so with a smaller sampling budget, setting a new state-of-the-art result among sub-7B-parameter LLMs.
@@
Table 2: Comparison of Apollo accuracy against state-of-the-art models on the miniF2F-test dataset. Token budget is computed over all generated tokens by LLM.
Method Model size Sample budget Token budget miniF2F-test
@@
Hypertree Proof Search [26] 600M $64\times 5000$ - $41.0\%$
InternLM2.5-SP+BFS+CG [24] 7B $256\times 32\times 600$ - $65.9\%$
HunyuanProver v16+BFS+DC [42] 7B $600\times 8\times 400$ - $68.4\%$
BFS-Prover [25] 7B $2048\times 2\times 600$ - $70.8\%$
@@
DeepSeek-R1-Distill-Qwen-7B [43] 7B 32 - 42.6%
Leanabell-Prover-GD-RL [32] 7B $128$ - $61.1\%$
STP [33] 7B $25600$ - $67.6\%$
o3-mini [39] - 1 - $3.3\%$
$32$ - $24.6\%$
o4-mini [40] - 1 - $7.0\%$
Goedel-SFT [31] 7B 32 16.1K $57.6\%$
$3200$ 1.6M $62.7\%$
$25600$ 12.7M $64.7\%$
Kimina-Prover-Preview-Distill [34] 7B 1 4.4K $52.5\%$
$32$ 140K $63.1\%$
$1024$ 4.5M $70.8\%$
Goedel-V2 [41] 8B 32 174K $83.3\%$
$64$ 349K $84.0\%$
$128$ 699K $84.9\%$
@@
o3-mini + Apollo - 8 - 40.2% (+36.9%)
o4-mini + Apollo - 15 - 46.7% (+39.7%)
Goedel-Prover-SFT + Apollo 7B 362 179K 65.6% (+0.9%)
Kimina-Preview + Apollo 7B 307 1.3M 75.0% (+4.2%)
Goedel-V2 + Apollo 8B 63 344K 84.9% (+0.9%)
@@ 4.3 Distribution of Proof Lengths
Figure 4: Distribution of proof lengths for selected models before (left) and after (right) applying Apollo. In the “before” plots, proof lengths cover only proofs the base model proved without the help of Apollo; in the “after” plots, proof lengths cover only proofs proved with Apollo’s assistance.
@@ 4.4 Impact of recursion depth $r$ on proof accuracy
We evaluate how the recursion-depth parameter $r$ affects accuracy on MiniF2F-test. A recursion depth of $r=0$ corresponds to the base model’s standalone performance; higher $r$ values permit additional rounds of Apollo. Figure 5 plots accuracy versus $r$ for three models: o4-mini ($r=0\ldots 4$), Kimina-Prover-Preview-Distill-7B ($r=0\ldots 6$), and Goedel-Prover-SFT ($r=0\ldots 6$).
@@ 6 Conclusion
In this work, we present Apollo, a novel, modular pipeline that combines syntax cleaning, automatic solvers, and LLM-driven sub-proof generation to transform an LLM’s initial proof sketch into a fully verified Lean4 proof. This framework harnesses the full power of the Lean compiler and integrated solvers, merging them with LLM systems. Applied across five whole-proof generation models, ranging from general-purpose LLMs (o3-mini, o4-mini) to specialized provers (Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT, Goedel-V2), Apollo consistently establishes new best accuracies on the miniF2F benchmark while reducing sampling budgets by one to two orders of magnitude. Our empirical analysis shows that Apollo not only raises overall proof success rates but also guides models to generate longer, more structured proofs, as evidenced by a drastic increase in average successful proof length. We further demonstrate how the recursion-depth parameter $r$ trades off sample complexity against accuracy, achieving robust gains with only a few hundred samples. We believe that Apollo’s collaboration between LLMs, automatic solvers, and the Lean compiler defines a paradigm of agentic systems that produce high-quality, verifiable proofs for increasingly challenging theorems.
Appendix A Additional implementation details
Algorithm 1 presents a high-level pseudo code of the Apollo framework. We highlight that the framework is built-upon the recursive function calls up to a maximum recursion depth given by parameter $r$. $\mathbf{LLM}$ refers to a proof generator model that takes in a proof statement and outputs a formal proof attempt. $\mathbf{LeanCompiler}$ refers to Lean REPL that verifies the proof and outputs True or False. In the case of False, it also provides a detailed list of errors, their locations and types along with the error messages generated by the Lean compiler. Other functions in the pseudo code ($\mathbf{SyntaxRefiner,Sorrifier,AutoSolver,ProofAssembler}$) represent the sub-modules described in the main text. To give deeper insight into Apollo’s design, we now outline the rationale and responsibilities of each core module.
A.1 SyntaxRefiner
The SyntaxRefiner class applies a set of rule-based corrections to eliminate syntax errors in LLM-generated Lean code. During model experimentation, we collected a corpus of common mistakes, such as stray Lean 3 keywords (from, beginend), misplaced tactics, or missing delimiters, and encoded each repair as a regular-expression rule. For example:
  • Convert [from by] to [:= by].
  • Replace Lean 3 blocks (beginend) with the corresponding Lean 4 structure.
  • Ensure that commands like rw use correctly placed square brackets to avoid compilation errors.
In more complex cases, multiple regex patterns and replacements are composed to restore valid Lean 4 syntax without altering the logical content of the proof sketch.
A.2 Sorrifier
The Sorrifier module uses a feedback loop between Apollo and the Lean REPL as follows:
  1. 1.
    Apollo sends the current proof candidate to REPL.
  2. 2.
    The REPL returns a list of error messages and their source locations.
  3. 3.
    Apollo parses the proof into a syntax tree, navigates to each failing sub-lemma, and attempts to fix it or replace it with a sorry placeholder.
  4. 4.
    The updated proof is re-submitted to REPL, and this cycle repeats until no error remain.
If the REPL indicates the original theorem statement itself is malformed (e.g. due to LLM hallucination), Apollo abandons the current proof and requests a fresh generation from LLM. One example of malformed formal statement is if LLM introduces a variable in a hypothesis that was not previously defined. Such case would trigger a compilation error in the formal statement; therefore, when parsing the proof, Apollo will catch this error and request a fresh proof from LLM. To guarantee consistent variable types during lemma extraction and proof assembly, Apollo sets the following Lean options to true:
1set_option pp.instanceTypes true
2set_option pp.numericTypes true
3set_option pp.coercions.types true
4set_option pp.letVarTypes true
5set_option pp.structureInstanceTypes true
6set_option pp.instanceTypes true
7set_option pp.mvars.withType true
8set_option pp.coercions true
9set_option pp.funBinderTypes true
10set_option pp.piBinderTypes true
so that the compiler reports explicit types (e.g. $\mathbb{Z}$ vs. $\mathbb{R}$). This fail-safe prevents trivialization of theorems and ensures correct typing in the final assembled proof.
A.3 AutoSolver
The AutoSolver module takes the “sorrified” proof and applies a sequence of tactics to close as many goals as possible:
  1. 1.
    Hint-based tactics. It first invokes hint, which proposes candidate tactics. We filter out any suggestions that merely progress the goal and retain only those that fully discharge it.
  2. 2.
    Built-in tactics. If hint fails, AutoSolver systematically tries a suite of powerful Lean tactics, such as nlinarith, norm_num, norm_cast, ring_nf, simp, simp_all, and others, often combining them with try to catch failures without aborting.
  3. 3.
    Recursive fallback. Any goals that remain unsolved are reported back to Apollo. Apollo queries the REPL for the exact goal statements, spawns corresponding sub-lemmas, and then invokes the full repair loop (LLM generation, SyntaxRefiner, Sorrifier, and AutoSolver) on each sub-lemma.
This tiered strategy maximizes automation while ensuring that challenging subgoals are handled via targeted recursion.
Algorithm 1 Apollo Pseudocode for Formal Theorem Proving
1:problem statement psps, current recursion depth $r_{current}$, maximum depth $r$
2:if $r_{current}$ is not initialized then
3:  $r_{current}\leftarrow 0$
4:end if
5:if $r_{current}>r$ then
6:  return sorry $\triangleright$ If recursion depth reached, return sorry
7:end if
[…]
12:if $\mathbf{LeanCompiler}(proof)$ == True then
13:  return $\color[rgb]{0,.5,.5}proof$
14:end if
[…]
18:for $i=1$ to $n$ do
19:  $ps_{goal}\leftarrow\mathbf{GetTheoremGoal}(proof,i)$
20:  $sub\_ps\leftarrow\mathbf{TransformGoal}(ps_{goal})$
21:  $sub\_proofs[i]\leftarrow\mathbf{Apollo}(sub\_ps,r_{current},r)$
22:end for
[…]
24:return repaired_proofrepaired\_proof
Appendix B An example of Apollo proof repair procedure
Figure 6 shows a detailed run of the Apollo framework on mathd_algebra_332. First, the LLM generates an initial proof sketch, which fails to type-check in Lean. Apollo then applies its Syntax Refiner to correct simple errors (e.g. changing “from by” to “:= by”) and to strip out all comment blocks. Next, the Sorrifier replaces each failing sub-lemma with a sorry statement (six in this example). The AutoSolver module then attempts to discharge these sub-lemmas: it resolves four of them, leaving lemmas #5 and #6 unsolved. Apollo recurses on those two: each is passed again through LLM, Syntax Refiner, Sorrifier, and AutoSolver. After this single recursive iteration, all goals are closed. Finally, Apollo reassembles the full proof and returns it to the user.
This example illustrates Apollo’s repair logic. In this case the proof was fixed in one iteration; more complex theorems may require deeper recursive repair.
Figure 6: Step-by-step Apollo repair on mathd_algebra_332. The initial proof and all sub-lemma proofs are generated with o4-mini, then systematically repaired and reassembled by Apollo.
Figure 7: Proof attempt of mathd_algebra_158 produced by the base model (o4-mini) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Appendix C Comparison between base proofs generated from LLMs and Apollo-repaired counterparts
In this section, we present and discuss examples of initial proof attempts alongside their Apollo-repaired counterparts.
Figures 7, 8, 9 illustrate cases where Apollo repairs the proof without invoking the LLM again. In each example, the LLM struggles to generate the precise tactics needed to close certain goals; however, we find that regenerating the entire proof is unnecessary. Instead, built-in Lean solvers can discharge those subgoals directly. Moreover, in Figure 7, problematic block #3 is resolved simply by removing it, since the surrounding proof context is correct. Thus, omitting problematic lines can sometimes yield a valid proof.
In Figure 10, one goal is expanded via an LLM query, but the model injects extra tactics that trigger a no goals to be solved error. Apollo then repairs the first block using a combination of LLM generation and AutoSolver, while the second block is removed entirely.
Figure 11 illustrates a case where the model fails to apply nlinarith to discharge the goal. We observe that, when uncertain, the model often resorts to broad tactics in hopes of closing the goal. Here, the goal is too complex for nlinarith, so Apollo leverages both the LLM and AutoSolver to guide the proof to completion.
Figure 12 illustrates an example of proof that required a recursion depth $r=5$ to repair the initial proof attempt. We observe that LLM’s proof structure is correct; however, in many cases it over-relies on built-in solvers and rewrites. However, since the goals are too complex, this approach leads to a total of nine error blocks. Repairing those blocks requires aid from both AutoSolver module of Apollo and LLM itself. We show that if LLM grasps the correct approach, then Apollo is able to repair fine grained logic errors to produce a correct proof.
Figure 13 illustrates a case where the LLM produces an incomplete proof sketch. The first half of the proof is correct, but the model fails to discharge the remaining goal and instead uses all_goals positivity. Apollo detects this error and performs two additional recursive repair iterations to complete the proof.
Figure 14 illustrates a case where the base model takes a lazy approach by trying to close all goals with linarith. In contrast, Apollo’s repaired proof performs additional setup and discharges each goal by first proving the necessary auxiliary hypotheses. This example shows that, although the model often understands the high-level strategy, it lacks the fine-grained tactics (and compiler feedback) needed to close subgoals. Apollo decomposes the proof, identifies the failure points, and applies targeted repairs to generate a fully verified solution.
Figure 15 shows another scenario in which Apollo successfully closes the first two blocks with linarith, but the final block requires a deeper reasoning chain. Apollo augments the proof by introducing and proving the missing lemmas, which leads to a correct solution with a series of rewrites.
Figure 16 presents an example of a very large repair. As in Figure 12, Apollo requires $r=5$ to fully fix the proof. The final proof length increases from 100 lines to 216, which means Apollo applied roughly $\times 2.2$ more tactics to close the goal. The repair relied on combined effort from AutoSolver and the LLM to close all goals. We show that, even for large, complex proofs, Apollo can repair each failing sub-block and generate a correct solution.
Figure 8: Proof attempt of mathd_algebra_141 produced by the base model (o4-mini) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 9: Proof attempt of imo_1983_p6 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 10: Proof attempt of induction_pord1p1on2powklt5on2 produced by the base model (Goedel-Prover-SFT) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 11: Proof attempt of mathd_algebra_293 produced by the base model (Goedel-Prover-SFT) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 12: Proof attempt of mathd_algebra_184 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 13: Proof attempt of mathd_algebra_293 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 14: Proof attempt of algebra_2varlineareq_fp3zeq11_3tfm1m5zeqn68_feqn10_zeq7 produced by the base model (o4-mini) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 15: Proof attempt of amc12a_2019_p12 produced by the base model (Goedel-Prover-SFT) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 16: Proof attempt of mathd_algebra_289 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals. Import headers are the same as other presented examples, for visibility they were omitted in this diagram.